OpenStudy (anonymous):

. In a city council election, 2,445 votes were cast for one of the candidates. If the candidate received at least one-third of the votes, how many voters voted in the election? a. Write an inequality to represent this situation, using x to represent the number of voters that voted in the election.

OpenStudy (help!!!!):

Yeah so do you have an idea how to make an inequality?

OpenStudy (anonymous):

yes

OpenStudy (help!!!!):

how would you write it? Give it a shot

OpenStudy (anonymous):

2445x<1/3

OpenStudy (help!!!!):

\[\frac{ 1 }{ 3 }x =\]\[\frac{ 1 }{ 3 }x=2445\]

OpenStudy (anonymous):

so it 1/3<2445

OpenStudy (anonymous):

right

OpenStudy (help!!!!):

add x

OpenStudy (help!!!!):

1/3x=2445

OpenStudy (anonymous):

how do i solve it b. Solve the inequality by isolating the variable. Show all work leading to your answer.

OpenStudy (help!!!!):

\[\frac{ 1 }{3 }x \le2445\]

OpenStudy (help!!!!):